newtons law of cooling states that the rate at which an object cool is directly proportional to the difference in the temperature between the object and its surrounding medium . Newtons law can be used to show that under certain conditions T (in degree Celcius) of an obeject at t (in hours) is given T=75 e^(-2t) express t is a function of T
Answer by Boreal(15235)

T=75 e^(-2t) ln of both sides
ln (T)=ln75-2t
2t=ln(75)-ln(T)
t=(1/2)[ln(75)-ln(T)]
t=(1/2)[ln(75/T)], and there are other ways to write it.